40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 11 15:04:52 2009, Alberto, Update, ABSL, Working on the AP table 
    Reply  Wed Nov 11 16:18:57 2009, Alberto, Update, ABSL, Working on the AP table 
       Reply  Wed Nov 11 17:10:51 2009, Jenne, Update, ABSL, Working on the AP table 
       Reply  Wed Nov 11 17:33:54 2009, Koji, Update, ABSL, Working on the AP table 
          Reply  Thu Nov 12 11:34:38 2009, Alberto, Update, ABSL, Working on the AP table 
             Reply  Thu Nov 12 16:53:59 2009, Alberto, Update, ABSL, Working on the AP table 
Message ID: 2239     Entry time: Wed Nov 11 16:18:57 2009     In reply to: 2238     Reply to this: 2240   2241
Author: Alberto 
Type: Update 
Category: ABSL 
Subject: Working on the AP table 

Quote:

I've opened the AP table and I'm working on it.

I re-aligned the Faraday on the AP table. I also aligned the beam to the periscope on the PSL and all the other optics along the beam path.  Now I have a nice NPRO beam at the PLL which overlaps with the PSL beam. The alignment has to be further improved because I see no beat yet.

I wonder if the all the tinkering on the PSL laser done recently to revive the power has changed the PSL NPRO temperature and so its frequency. That could also explain why the beat doesn't show up at the same temperature of the NPRO as I used to operate it. Although I scanned the NPRO temperature +/- 2 deg and didn't see the beat. So maybe the misalignment is the casue.

Not feeling very well right now. I need to go home for a while.

AP table closed at the moment.

NPRO shutter closed

ELOG V3.1.3-